A shop owner has 40 pounds of cashews worth $8 per pound and peanuts worth $2 per pound. How many pounds of peanuts should the owner mix with the 40 pounds of cashews to make a mixture worth $4 per pound?
Tried using a system of equations, but need some help. Thank you.

Answer by checkley79(3341)

8*40+2X=4(40+X)
320+2X=160+4X
2X-4X=160-320
-2X=-160
X=-160/-2
X=80 POUNDS OF PEANUTS ARE NEEDED.
PROOF:
320+2*80=160+4*80
320+160=160+320
480=480